# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### v3

@@ 6.1 Learning Performances
The learning performances for all the three cases are reported in Figure 10 and their numerical values tabulated in Table 2. The Figure shows the mean control or assimilation performances (first or second column, respectively) with a dashed line, and a confidence interval with a standard deviation above and below the mean performances. The proposed RT paradigm learns to control effectively the system in a limited number of iterations for all test cases, showing good sample complexity. Indeed, the optimized policy performs up to ten times better than the initial one, as shown by the average control performance ($\mathcal{L}_{a}$). The large standard deviation at the beginning of the training of T2 and T3 can be explained by the initial exploration phase of the model-free loop. However, the standard deviation gets roughly constant as the policy optimization converges to a specific region of the search space, as shown in Figures 9(c) and 9(d). Analysing the adjoint-based assimilation performances, Figures 9(b), 9(d) and 9(f) show that the methodology is able to tailor the digital twin on the current system specifics within five to ten iterations. At the end of the training process, the assimilation reduces the initial cost $\mathcal{L}^{(0)}_{p}$ from the initial guess $\bm{w}_{p}$ by a factor $10$, $100$ and $10000$ for T1, T2 and T3, respectively.
This, in turn, reflects on the model-based policy optimization (Step 5). In fact, as the digital twin predictions of the system response improve, the quality of the policy optimized “offline” increases. This also affects the policy switch (Step 6), since the performances on the twin become more representative of the real responses. The effect of Step 6 is illustrated in Fig. 11. All experiments begin with the model-free policy driving the interactions with the real environment, i.e. as “live policy”, while the model-based is initially set as “idle”. As shown in Figure 11, the switch between model-free and model-based policies occurs at $n_{e}=7$ in T1 and $n_{e}=4$ for both T2 and T3, as $C_{\pi}>P_{\pi}$ in the policy switch in Figure 3. After the switch, the model-free control performance registers a steeper improvement despite remaining idle. This is due to a combination of factors: (1) the model-based acts as an “expert” for the model-free, providing high quality samples recorded in $\mathcal{D}_{\mathcal{L}}$ from which the model-free loop can learn; and (2) the critic function approximation gets better at predicting the value associated to the recorded transitions. Interestingly, in both T1 and T3 (Figure 10(a), 10(c)), the model-free learns a better policy by imitating the model-based counterpart, eventually “surpassing the master” and earning again the status of a “live” policy again within a few episodes.
@@
Figure 11: Policy switch detail. Switch between “live” and “idle” policies, during interactions with real system.
@@ 6.2 Performances on Wind Turbine Control
The digital twin performances on T1 are shown in Figure 12. Figure 11(a) shows the average prediction versus the average system evolution as a function of time, while Figure 11(b) plots one versus the other. We conclude that the digital twin successfully tracks the dynamics of the real system albeit with a systematic under-estimation of roughly 1 (rad /s) or 9.55 (RPM), especially at low rotational speeds. Whether this bias is inherently linked with the chosen parametrization or an insufficient number of samples requires further investigation. Differently from T2 and T3, the coefficients to be identified are unknown a priori, as the initial coefficients $\bm{w}_{p}^{0}$, derived in the work of Saint-Drenan et al. (2020), describe a power curve in steady wind conditions. Therefore, the average assimilation loss $\mathcal{L}_{p}\approx\mathcal{O}(10^{-1})$ highlights the capability of the algorithm to adapt online, possibly taking into consideration changes in the operative conditions of the system itself.
Figure 12: Digital Twin agreement with real system. Evaluation of the optimal $\bm{w}_{p}=\bm{w}_{p}^{*}$ with randomized initial conditions and disturbances. Time-series (a) and scatter plot (b), plotting real state on the x-axis and predicted one on the y-axis.
The control performances are summarized in Figure 13. The controller is able to keep the system close to the desired set-points, as shown in Figures 12(a) and 12(b). This is especially true for the power tracking, always kept very close to $\tilde{P}$ despite the noticeably varying wind speeds (Figure 12(b)). On the other hand, the controller consistently keeps the generator speed ($\omega_{g}$) below-rated conditions by approximately 8 rad/s. This lower-than-rated rotational speed reflects on the actuated torque, in Figure 12(d), which is always in the middle between the allowed action space. In addition, the very fast actuation rate on the torque ($\dot{\tau}$ = 15 kN/s) allows to rapidly adapt in case of abrupt changes in the wind speed, while fine tuning the actuation with the pitch angle $\beta$. The latter follows consistently the wind speed, which is desirable. At higher wind speeds than rated ($u_{\infty}>11.4$ m/s) it consistently plays on the blades’ angle to reduce the aerodynamic torque to keep the system close to desired conditions. This aerodynamic de-rating of the rotor is visible in Figure 14(b) which shows a peak of $C_{p}$ at rated conditions ($u_{\infty}\approx 11$ (m/s)) with a $C_{p}=0.42$, steadily decreasing for higher winds.
@@
Figure 13: Control performances. System evolution against set-points (dashed red line) in (a) and (b), filtered actuations in (c) and (d).
Figure 14: Power coefficient $C_{p}$ against operative wind speeds.
The model-based and model-free policies are compared in Figure 15. The two optimization loops converge to essentially the same parametrizations, leading to the same policy functions, as also supported by the very similar mean cost functional $\mathcal{L}_{a}$ listed in Table 2. The only (minor) difference appears in the threshold of torque actuation. In both cases, when the error on the rotational speed goes negative ($e_{i}=\omega_{g}-\tilde{\omega}_{g}<0$), the controller reduces the pitch angle (Figures 14(a),14(c)) to increase the aerodynamic efficiency and, consequently, the aerodynamic torque, with the goal of increasing the rotational speed to get closer to the set point. The opposite is true when the generator is rotating too fast: the controller increases the angle of attack of the blades to slow the generator down. The integral of the error over the past revolution, $\int e$, essentially decides on the slopes of the control surface. Interestingly, looking at the torque actuation shown in Figures 14(b) and 14(d), it seems that this control actuation is almost insensitive to this control input, relying on the scalar error solely. This control surface is directly connected with the objective of keeping the power around the rated value of $P=5$ MW. In fact, as the error becomes negative, the controller increases the torque to keep the power at the end of the generator constant ($P_{g}=\omega_{g}\cdot\tau_{g}$). On the other hand, when the error increases, it reduces the resistive torque, relying on the pitch actuator to reduce the speed while keeping the power constant.
@@
Figure 15: Normalized policy contours, for the model-based (firsts two) and model-free (following) actor. As a function of the error $e_{k}(t)=\omega_{g}(t)-\tilde{\omega}_{g}(t)$ and its integral over the past revolution.
We conclude this analysis by investigating the robustness of the identified models and the control laws. Figure 15(b) shows the distribution of the average (a) control cost function for both model based and model free and (b) assimilation, obtained by running one-hundred simulations on randomized initial conditions and wind profiles (unseen during the training phase) and using the identified model parameters $\bm{w}_{p}^{*}$ and policy parameters $\bm{w}_{a,mb}^{*}$, $\bm{w}_{a,mf}^{*}$.
The control performances shows a multi-modal distribution in Figure 15(b), with the strongest peak in the region of better performances ( $\mathcal{L}_{a}<\mathcal{O}(10^{2})$). This shows that at least one local minima exists within the range of investigate conditions and the training can eventually fall into the sub-optimal one depending on the combination of initial condition and wind profiles. A possible solution to mitigate the sensitivity to the training condition could be to extend the training to a broader range of wind profiles and/or initial conditions.
The assimilations performances, on the other hand, are reliably on-track with respect to what experienced during the learning phase. The digital twin tracks the real system with very small deviation ($\mathcal{L}_{p}<\mathcal{O}(0.5\cdot 10^{5})$) for most experienced operative conditions.
Figure 16: Robustness analysis of the assimilated closure (a) and control policies (b)
Figure 16: Robustness analysis of the assimilated closure (a) and control policies (b)
@@ 6.3 Performances on a FWMA Trajectory Control
Figure 17: Comparison of the time evolution of the drone position for the real environment (solid lines) and the digital twin (dashed line). The right plot shows the high correlation between the norm of the predicted state $\bm{s}=[\dot{x},\dot{y},x,y]$ and the real state $\bm{\check{s}}$ for an episode length.
@@
Figure 18 characterizes the identified optimal policy, depicting state variables (17(a) and 17(b)) and control actions (17(c)). The figures evidence that the drone managed to closely reach its target ($\bm{x}_{f}=5$ m and $\bm{\dot{x}}_{f}=0$, as fast as possible) thanks to the control of the stroke plane angle $\beta_{w}$ and the flapping amplitude $A_{\phi}$. The control of the latter is mainly formed of three phases in a kind of bang-bang fashion. First, the controller sets the drone’s wings at their maximum angle of attack ($A_{\phi}=88$ deg) to reach the target position as fast as possible. Then, it decelerates and stabilizes the drone picking a flapping angle close to the minimal one ($55$ deg), with a minimal overshoot. Finally, the wings’ angle of attack is increased again, to balance the body weight with the aerodynamic lift. This force is also slightly modulated by the other control variable ($\beta_{w}$) which mainly drives the horizontal motion of the drone. In the same way, the controller pitches the stroke plane up to the maximum angle ($\beta_{w}=-30$ deg) and then stabilizes it around an angle that allows it to stay close to the desired position, $x_{f}$, despite the headwind. The influence of the gust is visible on $\beta_{w}$ even though the wind effect is still very limited, due to the selected mass and surface of the body.
@@
Figure 19 analyses the policies found by the model-free (Figures 18(a) - 18(d)) and model-based (Figures 18(e) - 18(h)) loops. To this end, we fix two of the four inputs to zero and span on the other two dimensions, recording the corresponding action. The two policies are remarkably similar, converging to essentially the same parametrization. The main difference between the two lies in the concavity of the hyperbolic tangent, as dictated by the biases. This affects the localization of the switching between the two bounds of the actuation space, possibly explaining the differences in their associated costs presented in Table 2. In both cases, for what concerns $\hat{A}_{\phi}$, the border between the two bounds lies on the diagonal that goes from $(\hat{\dot{e}}_{x},\hat{e}_{x})=(-1.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x},)=(0.25,1.0)$, see Figure 18(a). The same trend is observed in Figure 18(b), which analyses the dependencies from $e_{z}$ and its time derivative. In this case, however, the upper point of the diagonal lies around $(\hat{\dot{e}}_{z},\hat{e}_{z})=(0.0,1.0)$, thus increasing $\hat{A}_{\phi}$ as the rate of change of $e_{z}$ goes negative, and decreasing it otherwise. Interestingly, the model-based policy here shows a broader decision zone, as depicted in Figure 9(f). Now, analysing the actuation on $\beta_{w}$, the threshold appears to be positioned in the diagonals $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.75,1.0)$. The sensitivity to the $z-$errors instead show it to lie in $(\hat{\dot{e}}_{z},\hat{e}_{z})=(-1.0,0.0)$ and $(\hat{\dot{e}}_{z},\hat{e}_{z})=(1,0.5)$. Therefore, it seems that the time variation of the rate of change is the most important input on the $x-$dimension, whereas in the $z-$coordinate, it is the scalar $\hat{e}_{z}$. Arguably this may be caused by the headwind, oriented as opposite of the $x-$axis. Overall, there appears to be two main mechanisms at play: (1) the errors and their rate of change seems to have the same importance when deciding how to control on $\hat{A}_{\phi}$ being slightly more sensitive to the scalar instantaneous errors; and (2) when deciding on how to actuate $\beta_{w}$, the controller balances the scalar instantaneous error in one dimension and its rate of change in the other. In fact, when $\dot{e}_{x}$ increases, it actuates on $\beta_{w}$ to get back to nominal conditions, and the same is observed on analysing the reactivity to the instantaneous error $e_{z}$ and $e_{x}$.
@@
Figure 18: Time evolution of the positions $x$ and $y$ (a) and the velocities $\dot{x}$ and $\dot{y}$ (b) of the flapping wing micro air vehicle driven by the control actions $\bm{a}=[A_{\phi},\beta_{w}]$ (c). The optimal control policy manages to lead the drone to its target $(x,y)=(5,5)$ and $(\dot{x},\dot{y})=(0,0)$ within one episode time with minor oscillations that results from the gust.
@@
Figure 19: Comparison of model-based and model-free policies.
@@
Figure 20: Violin plots of the model-free and model-based policies compute over 100 episodes. Both policies result in similar distributions with a lower average cost for the model-free policy

### v4

@@ 6.1 Learning Performances
The learning performances for all the three cases are reported in Figure 10 and their numerical values tabulated in Table 2. The Figure shows the mean control or assimilation performances (first or second column, respectively) with a dashed line, and a confidence interval with a standard deviation above and below the mean performances. The proposed RT algorithm learns to control effectively the system in a limited number of iterations for all test cases, showing good sample efficiency. Indeed, the optimized policy performs up to ten times better than the initial one, as shown by the average control performance ($\mathcal{L}_{a}$). The large standard deviation at the beginning of the training of T2 and T3 can be explained by the initial exploration phase of the model-free loop. However, the standard deviation gets roughly constant as the policy optimization converges to a specific region of the search space, as shown in Figures 10(c) and 10(d). Analysing the adjoint-based assimilation performances, Figures 10(b), 10(d) and 10(f) show that the methodology is able to tailor the digital twin on the current system specifics within five to ten iterations. At the end of the training process, the assimilation reduces the initial cost $\mathcal{L}^{(0)}_{p}$ from the initial guess $\bm{w}_{p}$ by a factor $10$, $100$ and $10000$ for T1, T2 and T3, respectively.
This, in turn, reflects on the model-based policy optimization (Step 5). In fact, as the digital twin predictions of the system response improve, the quality of the policy optimized “offline” increases. This also affects the policy switch (Step 6), since the performances on the twin become more representative of the real responses. The effect of Step 6 is illustrated in Fig. 11. All experiments begin with the model-free policy driving the interactions with the real environment, i.e. as “live policy”, while the model-based is initially set as “idle”. As shown in Figure 11, the switch between model-free and model-based policies occurs at $n_{e}=7$ in T1 and $n_{e}=4$ for both T2 and T3, as $C_{\pi}>P_{\pi}$ in the policy switch in Figure 3. After the switch, the model-free control performance registers a steeper improvement despite remaining idle. This is due to a combination of factors: (1) the model-based acts as an “expert” for the model-free, providing high quality samples recorded in $\mathcal{D}_{\mathcal{L}}$ from which the model-free loop can learn; and (2) the critic function approximation gets better at predicting the value associated to the recorded transitions. Interestingly, in both T1 and T3 (Figure 11(a), 11(c)), the model-free learns a better policy by imitating the model-based counterpart, eventually “surpassing the master” and earning again the status of a “live” policy again within a few episodes.
@@
Figure 11: Policy switch detail. Switch between “live” and “idle” policies, during interactions with the real system.
@@ 6.2 Performances on Wind Turbine Control
The digital twin performances on T1 are shown in Figure 12. Figure 12(a) shows the average prediction versus the average system evolution as a function of time, while Figure 12(b) plots one versus the other. We conclude that the digital twin successfully tracks the dynamics of the real system albeit with a systematic under-estimation of roughly 1 (rad /s) or 9.55 (RPM), especially at low rotational speeds. Whether this bias is inherently linked with the chosen parametrization or an insufficient number of samples requires further investigation. Differently from T2 and T3, the coefficients to be identified are unknown a priori, as the initial coefficients $\bm{w}_{p}^{0}$, derived in the work of Saint-Drenan et al. [2020], describe a power curve in steady wind conditions. Therefore, the average assimilation loss $\mathcal{L}_{p}\approx\mathcal{O}(10^{-1})$ highlights the capability of the algorithm to adapt online, possibly taking into consideration changes in the operative conditions of the system itself.
Figure 12: T1 - Digital Twin agreement with real system. Evaluation of the optimal $\bm{w}_{p}=\bm{w}_{p}^{*}$ with randomized initial conditions and disturbances. Time-series (a) and scatter plot (b), plotting real state on the x-axis and predicted one on the y-axis.
The control performances are summarized in Figure 13. The controller is able to keep the system close to the desired set-points, as shown in Figures 13(a) and 13(b). This is especially true for the power tracking, always kept very close to $\tilde{P}$ despite the noticeably varying wind speeds (Figure 13(b)). On the other hand, the controller consistently keeps the generator speed ($\omega_{g}$) below-rated conditions by approximately 8 rad/s. This lower-than-rated rotational speed reflects on the actuated torque, in Figure 13(d), which is always in the middle between the allowed action space. In addition, the very fast actuation rate on the torque ($\dot{\tau}$ = 15 kN/s) allows to rapidly adapt in case of abrupt changes in the wind speed, while fine tuning the actuation with the pitch angle $\beta$. The latter follows consistently the wind speed, which is desirable. At higher wind speeds than rated ($u_{\infty}>11.4$ m/s) it consistently plays on the blades’ angle to reduce the aerodynamic torque to keep the system close to desired conditions. This aerodynamic de-rating of the rotor is visible in Figure 15(b) which shows a peak of $C_{p}$ at rated conditions ($u_{\infty}\approx 11$ (m/s)) with a $C_{p}=0.42$, steadily decreasing for higher winds.
@@
Figure 13: T1 - Control performances. System evolution against set-points (dashed red line) in (a) and (b), filtered actuation in (c) and (d).
Figure 14: T1 - Power coefficient ($C_{p}$) against operative wind speeds.
The model-based and model-free policies are compared in Figure 15. The two optimization loops converge to essentially the same parametrizations, leading to the same policy functions, as also supported by the very similar mean cost functional $\mathcal{L}_{a}$ listed in Table 2. The only (minor) difference appears in the threshold of torque actuation. In both cases, when the error on the rotational speed goes negative ($e_{i}=\omega_{g}-\tilde{\omega}_{g}<0$), the controller reduces the pitch angle (Figures 15(a),15(c)) to increase the aerodynamic efficiency and, consequently, the aerodynamic torque, with the goal of increasing the rotational speed to get closer to the set point. The opposite is true when the generator is rotating too fast: the controller increases the angle of attack of the blades to slow the generator down. The integral of the error over the past revolution, $\int e$, essentially decides on the slopes of the control surface. Interestingly, looking at the torque actuation shown in Figures 15(b) and 15(d), it seems that this control actuation is almost insensitive to this control input, relying on the scalar error solely. This control surface is directly connected with the objective of keeping the power around the rated value of $P=5$ MW. In fact, as the error becomes negative, the controller increases the torque to keep the power at the end of the generator constant ($P_{g}=\omega_{g}\cdot\tau_{g}$). On the other hand, when the error increases, it reduces the resistive torque, relying on the pitch actuator to reduce the speed while keeping the power constant.
@@
Figure 15: T1 - Normalized policy contours. Model-based (a - b) and model-free (c - d) actors, as a function of the error $e_{k}(t)=\omega_{g}(t)-\tilde{\omega}_{g}(t)$ and its integral over the past revolution.
We conclude this analysis by investigating the robustness of the identified models and the control laws. Figure 16 shows the distribution of the average (a) control cost function for both model based and model free and (b) assimilation, obtained by running one-hundred simulations on randomized initial conditions and wind profiles (unseen during the training phase) and using the identified model parameters $\bm{w}_{p}^{*}$ and policy parameters $\bm{w}_{a,mb}^{*}$, $\bm{w}_{a,mf}^{*}$. The control performances shows a multi-modal distribution in Figure 16(b), with the strongest peak in the region of better performances ( $\mathcal{L}_{a}<\mathcal{O}(10^{2})$). This shows that at least one local minima exists within the range of investigate conditions and the training can eventually fall into the sub-optimal one depending on the combination of initial condition and wind profiles. A possible solution to mitigate the sensitivity to the training condition could be to extend the training to a broader range of wind profiles and/or initial conditions.
The assimilations performances, on the other hand, are reliably on-track with respect to what experienced during the learning phase. The digital twin tracks the real system with very small deviation ($\mathcal{L}_{p}<\mathcal{O}(10^{2})$) for most experienced operative conditions.
Figure 16: T1 - Robustness analysis of the control policies (a) and assimilated closure (b).
@@ 6.3 Performances on a FWMA Trajectory Control
Figure 17: T2 - Comparison of the time evolution of the drone position for the real environment (solid lines) and the digital twin (dashed line). The right plot shows the high correlation between the norm of the predicted state $\bm{s}=[\dot{x},\dot{y},x,y]$ and the real state $\bm{\check{s}}$ for an episode length.
@@
Figure 18 characterizes the identified optimal policy, depicting state variables (18(a) and 18(b)) and control actions (18(c)). The figures evidence that the drone managed to closely reach its target ($\bm{x}_{f}=5$ m and $\bm{\dot{x}}_{f}=0$, as fast as possible) thanks to the control of the stroke plane angle $\beta_{w}$ and the flapping amplitude $A_{\phi}$. The control of the latter is mainly formed of three phases in a kind of bang-bang fashion. First, the controller sets the drone’s wings at their maximum angle of attack ($A_{\phi}=88$ deg) to reach the target position as fast as possible. Then, it decelerates and stabilizes the drone picking a flapping angle close to the minimal one ($55$ deg), with a minimal overshoot. Finally, the wings’ angle of attack is increased again, to balance the body weight with the aerodynamic lift. This force is also slightly modulated by the other control variable ($\beta_{w}$) which mainly drives the horizontal motion of the drone. In the same way, the controller pitches the stroke plane up to the maximum angle ($\beta_{w}=-30$ deg) and then stabilizes it around an angle that allows it to stay close to the desired position, $x_{f}$, despite the headwind. The influence of the gust is visible on $\beta_{w}$ even though the wind effect is still very limited, due to the selected mass and surface of the body.
@@
Figure 19 analyses the policies found by the model-free (Figures 19(a) - 19(d)) and model-based (Figures 19(e) - 19(h)) loops. To this end, we fix two of the four inputs to zero and span on the other two dimensions, recording the corresponding action. The two policies are remarkably similar, converging to essentially the same parametrization. The main difference between the two lies in the concavity of the hyperbolic tangent, as dictated by the biases. This affects the localization of the switching between the two bounds of the actuation space, possibly explaining the differences in their associated costs presented in Table 2. In both cases, for what concerns $\hat{A}_{\phi}$, the border between the two bounds lies on the diagonal that goes from $(\hat{\dot{e}}_{x},\hat{e}_{x})=(-1.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x},)=(0.25,1.0)$, see Figure 19(a). The same trend is observed in Figure 19(b), which analyses the dependencies from $e_{z}$ and its time derivative. In this case, however, the upper point of the diagonal lies around $(\hat{\dot{e}}_{z},\hat{e}_{z})=(0.0,1.0)$, thus increasing $\hat{A}_{\phi}$ as the rate of change of $e_{z}$ goes negative, and decreasing it otherwise. Interestingly, the model-based policy here shows a broader decision zone, as depicted in Figure 10(f). Now, analysing the actuation on $\beta_{w}$, the threshold appears to be positioned in the diagonals $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.75,1.0)$. The sensitivity to the $z-$errors instead show it to lie in $(\hat{\dot{e}}_{z},\hat{e}_{z})=(-1.0,0.0)$ and $(\hat{\dot{e}}_{z},\hat{e}_{z})=(1,0.5)$. Therefore, it seems that the time variation of the rate of change is the most important input on the $x-$dimension, whereas in the $z-$coordinate, it is the scalar $\hat{e}_{z}$. Arguably this may be caused by the headwind, oriented as opposite of the $x-$axis. Overall, there appears to be two main mechanisms at play: (1) the errors and their rate of change seems to have the same importance when deciding how to control on $\hat{A}_{\phi}$ being slightly more sensitive to the scalar instantaneous errors; and (2) when deciding on how to actuate $\beta_{w}$, the controller balances the scalar instantaneous error in one dimension and its rate of change in the other. In fact, when $\dot{e}_{x}$ increases, it actuates on $\beta_{w}$ to get back to nominal conditions, and the same is observed on analysing the reactivity to the instantaneous error $e_{z}$ and $e_{x}$.
@@
Figure 18: T2 - Time evolution of the positions $x$ and $y$ (a) and the velocities $\dot{x}$ and $\dot{y}$ (b) of the flapping wing micro air vehicle driven by the control actions $\bm{a}=[A_{\phi},\beta_{w}]$ (c). The optimal control policy manages to lead the drone to its target $(x,y)=(5,5)$ and $(\dot{x},\dot{y})=(0,0)$ within one episode time with minor oscillations that results from the gust.
@@
Figure 19: T2 - Comparison of model-based and model-free policies.
@@
Figure 20: T2 - Robustness analysis of model-free and model-based policies (a) and assimilated closure (b), compute over 100 episodes.